# Deep Optical Emission-Line Images of Nine Known and Three New Galactic Supernova Remnants

### v1

@@ 3.4 New Galactic SNRs
Following the identification by M. Drechsler of a large emission shell some eight degrees off the Galactic plane in the 1.4 GHz NRAO VLA Sky Survey (NVSS; Condon et al. 1998), exploratory images by B. Falls led to the discovery of faint optical emission matching the position of the faint radio shell. The remnant, roughly 1.8 degrees in diameter and centered at , was subsequently found to be mainly visible optically via [O iii] emission. Follow-up H and [O iii] images revealed a nearly complete emission shell brightest in the north and east. Nicknamed the ‘Atlas’ remnant due to the Atlas mountains of Morocco from where it was often imaged, Figure 35 shows our [O iii] (top) and H (bottom) images. In [O iii], the remnant exhibits a bright, flat northern rim, with its eastern edge unusually sharp and well defined. This sharp eastern edge is also brightest in the 1.4 GHz NVSS map. Although much of the western limb is faint and/or missing, there is considerable internal diffuse emission near the remnant center.
@@
Four exploratory long slit spectra taken along the nebula’s western limb using the MDM 2.4 telescope with the OSMOS spectrograph/camera found I(S ii]/I(H) ratio between 0.44 - 0.62 () confirming a shock nature. The [S ii] 6716/6731 ratio was also found to be between 1.30 and 1.45 suggesting electron densities at or near the low density limit, i.e., n = cm.
@@ 4 Discussion and Conclusions
As an example of the depth of our images, we show in Figure 40 a comparison of our H image with that of the VTSS on the G209.9-8.2 remnant. VTSS claims an H detection limit down to 2 Rayleighs in a 2 hr exposure using a 17 Å wide filter. Our H image is the product of 120 hr using 2 to 3 times wider filters with a resolution of 2.4, some 40 times better than the 98 listed for VTSS. While both images show similar emission features, our image with its substantially better angular resolution has a detection limit at least as deep as the VTSS. Moreover, remnant features can be more readily distinguished in our image from background emissions such as the emission along much of the western portion of the VTSS image arising from a bright and unrelated nearby H II region.
@@
G150.3+4.5: Our images of this poorly studied remnant represents the first optical investigation. As we have seen in other SNRs in our survey, deep [O iii] images proved invaluable in determining the remnant’s structure at a much high resolution than that of radio data. Our [O iii] images also unexpectantly uncovered a complete emission shell surrounding the variable X-ray binary source CI Cam which appears to have not been detected before.
@@
G321.3-3.9: This remnant displays an amazing amount of optical emission, and like that seen for some other remnants, it is particularly bright in [O iii]. It also displays two shell blowout features. While the remnant’s optical emission is in excellent agreement with the remnant’s brightest radio emission, it exhibits an enormous wealth of additional details unseen in the radio.
We also announce the discovery and present here deep H and [O iii] images of three new apparent Galactic remnants. Each of these new SNRs have extensive optical structures comparable in size and detail to that seen for the best studied and imaged known SNRs. For example, the extensive and beautiful [O iii] filaments of the large SNR G107.7-5.1 have few rivals in the other 300+ Galactic SNRs. The exceedingly faint but beautiful G209.9-8.2 remnant is one more example where [O iii] imaging reveals a SNR far better than deep H images. And the remarkable [O iii] ‘bubble’ in the G210.5+1.3 distinct from the its H emission filaments, best seen in the starless version presented here, gives a new prospective regarding the size of expansion blowouts.
@@
2) The great complexity of some remnants in terms of their H and [O iii] line emissions, seen in G70.0-21.5, G107.7-5.1, G119.5+10.5 (CTA 1), and G321.3-3.9, raises serious questions about the usefulness of spectral analysis results from just a handful of slit splectra taken on a few filaments. The [O iii] bubble of G210.5+1.3 dramatically shows that without deep emission line images besides H one may miss large portions of a remnant whose spectral properties differ greatly from other regions.
3) Remnant “blowout” features do not seem to be especially rare. In our survey of nine known and three new remnants, about 30% exhibited significant extended features outside their main shell; these include G70.0-21.5, G82.2+5.3, G210.5+1.3, and G321.3-3.9. Blowout features also appear common in radio data. A recent study of 36 SNRs using MeerKAT 1.3 GHz observations found nearly 50% of the remnants showed blowout or “ears” (Cotton et al., 2024). However, we caution that our small survey focused on particularly large remnants where their large size might be more sensitive to local interstellar density variations which subsequently show up as shock expansion like bubbles.

### v2

@@ 3.4 New Galactic SNRs
Following the identification by M. Drechsler of a large emission shell some eight degrees off the Galactic plane in the 1.4 GHz NRAO VLA Sky Survey (NVSS; Condon et al. 1998), exploratory images by B. Falls led to the discovery of faint optical emission matching the position of the faint radio shell. The remnant, roughly 1.8 degrees in diameter and centered at , was subsequently found to be mainly visible optically via [O III] emission. Follow-up H and [O III] images revealed a nearly complete emission shell brightest in the north and east. Nicknamed the ‘Atlas’ remnant due to the Atlas mountains of Morocco from where it was most often imaged, Figure 35 shows our [O III] (top) and H (bottom) images. In [O III], the remnant exhibits a bright, flat northern rim, with its eastern edge unusually sharp and well defined. This sharp eastern edge of [O III] emission matches exactly bright eastern emission arc seen in the 1.4 GHz NVSS map, whereas the much fainter western 1.4 GHz emission arc lies some farther west than the western edge of the detected [O III] emission. Although parts of the western limb is faint and/or missing in [O III], there is an unusual amount of internal diffuse emission near and north of the remnant center.
@@
Four exploratory long slit spectra taken along the nebula’s western limb using the MDM 2.4 telescope with the OSMOS spectrograph/camera found I(S II]/I(H) ratio between 0.44 - 0.62 () confirming a shock nature. The [S II] 6716/6731 ratio was also found to be between 1.30 and 1.45 suggesting electron densities at or near the low density limit, i.e., ne = cm-3.
@@ 4 Discussion and Conclusions
As an example of the depth of our images, we show in Figure 40 a comparison of our H image with that of the VTSS on the G209.9-8.2 remnant. VTSS claims an H detection limit down to 2 Rayleighs in a 2 hr exposure using a 17 Å wide filter. Our H image is the product of 120 hr using 2 to 3 times wider filters with a resolution of 2.4′′, some 40 times better than the 98′′ listed for VTSS. While both images show similar emission features, our image with its substantially better angular resolution has a detection limit at least as deep as the VTSS. Moreover, remnant features can be more readily distinguished in our image from background emissions such as the emission along much of the western portion of the VTSS image arising from a bright and unrelated nearby H II region.
@@
G150.3+4.5: Our images of this poorly studied remnant represents the first optical investigation. As we have seen in other SNRs in our survey, deep [O III] images proved invaluable in determining the remnant’s structure at a much high resolution than that of radio data. Our [O III] images also unexpectedly uncovered a complete emission shell surrounding the variable X-ray binary source CI Cam which appears to have not been detected before.
@@
G321.3-3.9: This remnant displays an amazing amount of optical emission, and like that seen for some other remnants, it is particularly bright in [O III]. It also displays two shell breakout features. While the remnant’s optical emission is in excellent agreement with the remnant’s brightest radio emission, it exhibits an enormous wealth of additional details unseen in the radio.
We also announce the discovery and present here deep H and [O III] images of three new apparent Galactic remnants. Each of these new SNRs have extensive optical structures comparable in size and detail to that seen for the best studied and imaged known SNRs. For example, the extensive and beautiful [O III] filaments of the large SNR G107.7-5.1 have few rivals in the other 300+ Galactic SNRs. The exceedingly faint but beautiful G209.9-8.2 remnant is one more example where [O III] imaging reveals a SNR far better than deep H images. And the remarkable [O III] ‘bubble’ in the G210.5+1.3 distinct from the its H emission filaments, best seen in the starless version presented here, gives a new prospective regarding the size of expansion breakouts.
@@
2) The great complexity of some remnants in terms of their H and [O III] line emissions, seen in G70.0-21.5, G107.7-5.1, G119.5+10.5 (CTA 1), and G321.3-3.9, raises serious questions about the usefulness of spectral analysis results from just a handful of slit spectra taken on a few filaments. The [O III] bubble of G210.5+1.3 dramatically shows that without deep emission line images besides H one may miss large portions of a remnant whose spectral properties differ greatly from other regions.
3) Remnant “breakout” features do not seem to be especially rare. In our survey of nine known and three new remnants, about 30% exhibited significant extended features outside their main shell; these include G70.0-21.5, G82.2+5.3, G210.5+1.3, and G321.3-3.9. breakout features also appear common in radio data. A recent study of 36 SNRs using MeerKAT 1.3 GHz observations found nearly 50% of the remnants showed breakout or “ears” (Cotton et al., 2024). However, we caution that our small survey focused on particularly large remnants where their large size might be more sensitive to local interstellar density variations which subsequently show up as shock expansion like bubbles.